An employee obtained a loan of P100,000 at the rate of 6% compounded in order to build a house. How much must he pay monthly to amortize the loan within a period of 10 years?


1
Expert's answer
2022-04-06T14:52:35-0400

"R = P \u00d7 r \/ [1 - (1 + r)^{-n}]"

"P=100000"

"r=0.06\/12=0.005"

"n=12\cdot 10=120"


"R=100000\cdot 0.005/(1-(1+0.005)^{-120})=P\ 1111.11"